OpenStudy (anonymous):

Evaluate these integrals using integration by parts? a) ∫ cos (In x) dx b) ∫ x tan^(2) x dx

OpenStudy (anonymous):

ln(x)=u

OpenStudy (anonymous):

du=1/x dx

OpenStudy (anonymous):

disregard my previous messages

OpenStudy (anonymous):

u=cos(ln(x)) dv= 1 du=-sin(ln(x))/x v=x

OpenStudy (anonymous):

I think this is a cyclic problem